OpenStudy (anonymous):

Find the value of Q in the following system so that the solution to the system is {(x,y):x-3y=4} x-3y=4 2x+Qy=8

OpenStudy (anonymous):

I think Q = -6

OpenStudy (anonymous):

But I think I didn't understand the question...

OpenStudy (anonymous):

You can write in this way : 2x-6y=8 2x+Qy=8 Now we have : 2x-6y=2x+Qy =>Q=_6 OK ?! Now answer friend !

OpenStudy (anonymous):

and now i'm sure...:D

OpenStudy (anonymous):

?? please explain!

OpenStudy (anonymous):

8/4 = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and this one? 2x / x = ?

OpenStudy (anonymous):

\[\frac{ 2x }{ x } = ?\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and Q / (-3y) should be 2 too, ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

OHHH i get it,.... but why all 2?

OpenStudy (anonymous):

ask this question from the designer of the question ... XD

OpenStudy (anonymous):

Because in all of the equetions we have + / * - for alllllllll

OpenStudy (anonymous):

He\she wanted this...

OpenStudy (anonymous):

D: so confusing!!

OpenStudy (anonymous):

:D , any problem?

OpenStudy (anonymous):

No problem ?:)

OpenStudy (anonymous):

@AkashdeepDeb , please look at our work...did we do that correct?

OpenStudy (anonymous):

is there any way to check it?

OpenStudy (anonymous):

Yes ! Finding X and Y too !!!

OpenStudy (anonymous):

finding X and Y in system? are you joking? :D

OpenStudy (anonymous):

gahhh ill be old :D

OpenStudy (anonymous):

@lavalava , how did you find this question?

OpenStudy (anonymous):

its part of my math subject -.- they dont explain very well

OpenStudy (anonymous):

nice...I think they NEVER explain:D I think we explained it correct :);)

OpenStudy (anonymous):

haha yes you did :)